Tony can paint a room in 4hr. Gina can paint he same room in 3 hr. Working together, how long will it take them to paint the room.
Answer by Earlsdon(6103)

Tony can paint the room in 4 hours, so this means he can paint 1/4 of the room in 1 hour.
Gina can paint the room in 3 hours, so this means she can paint 1/3 of the room in 1 hour.
Together, they can paint (1/4 + 1/3) = 7/12 of the room in 1 hour so it would take them 12/7 hours to paint the entire room working together.
12/7 hours = 1 hour and 25 minutes.